If a sample contains 6.75 mol of Na2SO4, how many moles of sodium atoms, sulfur atoms, and oxygen atoms does it have?


1
Expert's answer
2022-01-24T06:59:03-0500

2Na=23×2= 46g

S= 32g

4O= 16×4= 64g


6.75 mol of Na2SO4= 142× 6.75= 958.5g

Moles of sodium= 2moles

Moles of sulphur= 1.5moles

Moles of oxygen= 3.25moles